OpenStudy (anonymous):

Find the equation of the tangent line to the curve at the given point. y = 9x − 8sqrtx (1, 1)

OpenStudy (tkhunny):

Find the 1st Derivative. y' = ???

OpenStudy (anonymous):

If I did it right, I got 9-4x ^{-1/2}

OpenStudy (tkhunny):

Well, don't forget to write the entire equation or definition. It will keep you from becoming confused. y' = 9 - 4/sqrt{x} That is correct. Now, just evaluate it at x = 1 and you have your slope.

OpenStudy (anonymous):

m=5?

OpenStudy (tkhunny):

That's it. It's Point-Slope Time! Let's see the equation of the line.

OpenStudy (anonymous):

y-1=5(x-1) so, y=5x-4

OpenStudy (tkhunny):

Done. You're a pro.

OpenStudy (anonymous):

Thanks again! I think I was just over analyzing. I have literally worked out all 8 of my homework's and have made a list of questions just to put on here!